LIU-PSB Meeting 226 (Linac4-to-PSB LS2 schedule)

Europe/Zurich
864/2-B14 - SALLE J.B.ADAMS (CERN)

864/2-B14 - SALLE J.B.ADAMS

CERN

25
Show room on map
Gian Piero Di Giovanni (CERN)

Participants: Maria Elena Angoletta, Miroslav Georgiev Atanasov, Giulia Bellodi, Bartosz Przemyslaw Bielawski, Jean-François Comblin, Julie Coupard, James Devine, Gian Piero Di Giovanni, Eve Fortescue-Beck, Georgi Minchev Georgiev, Marine Gourber-Pace, Matthias Haase, Verena Kain, Bettina Mikulec, Gabor Petrika, Fabio Pozzi, Andrea Santamaria Garcia, Frank Schmidt.

Approval of minutes

The minutes of the LIU-PSB WG #224 and LIU-PSB WG #225 were approved.

Communications

General

  • The LIU event will be held from 13th to 15th February 2018, two nights, and will take place in Montreux. The discussion on the agenda is well advanced and the program is nearly finalised.

LIU-PT Meeting

  • There is an ongoing discussion about the internal restructuring of the LIU project, since the project is now entering a new phase with different demands.
  • The joint Linacs/LEIR/PSB/PS meetings are crucial to the project and will keep on taking place weekly during the installation phase, but not daily. Individual problems will be dealt with directly on site.
  • Concerning the start of LS2, the Linac2 access system has been dismantled and the PSB RF systems are currently being dismantled.

LIU Beam Performance Meeting

  • The last meeting was dedicated to an update on the brightness studies and transverse emittance blow-up at the PS injection.

LIU-PT meetings with dedicated equipment group

  • Final meetings with TE-MSC, EN-EL, EN-HE, and TE-ABT took place.

Open Actions

T. Dobers (2018-10-23): Produce the tilt measurement tool and perform measurements on spare / future magnets.

  • T. Dobers communicated by email that the tool was successfully tested to measure the roll (tilt) of the main dipoles of the Booster Ring. Action closed.

Linac-to-PSB LS2 schedule (J. Coupard)

J. Coupard presented the detailed schedule of the Linac4 connection. She reminded that S. Mataguez is the Activity Technical Coordinator and that she is the Facility Coordinator for the Linac4 connection.

Concerning the documentation, there are two ECRs missing, namely the LBE line and infrastructure upgrade for the Linac4 and the new BLM installation in the transfer lines. For the Work Package Analysis, all the work about the connection is in engineering check. The dedicated procedures, which concerns the equipment to be operated during the beam tests in 2019, are to be validated by everyone involved.

The schedule is divided between the work to be done in Linac4 and the connection interface in Linac2 and Linac4.

Completion of the Linac4 transfer line

  • The goal is to install the final configuration of the layout.
  • By the end of the year 2018, all equipment should be removed, so that the installation can be done in January 2018.
  • The Open Days have been explicitly included in the schedule, with one extra week without activities to perform a proper lock-out of the equipment.
  • There is an open question concerning the alignment of the Linac4 equipment since the tunnel is moving. Survey should give an assessment, and the vacuum will be re-opened if necessary.
  • The vacuum team will not be available later due to the high number of activities in other machines, so they will need to work first on the Linac4. Activities related to instrumentation, such as the wire scanners or SEM grids, will need to synchronize with the vacuum team.
  • The BLMs installation is included in May 2019.

Partial dismantling of Linac2

  • The removal of the Linac2 access system has been completed already, so the activities are on schedule. Currently, there is a cooldown period put in place by Radio Protection, and the dismantling will start after that. The beam stopper and beam dump will be removed.
  • Differential drawings exist and they should be released before the work starts.
  • Next year all the equipment will be disconnected, followed by the de-cabling campaign.

Reconfiguration of the power converters in B363

  • The present racks will be removed and installed in the new configuration with the new switchboard. It is not foreseen to happen before the 10th of December.
  • This will be followed by de-cabling, installation of the new configuration, then cabling.
  • The cooling and ventilation activities are programmed so that they will not clash with others.

Connection of the Linac4 to the LT.BHZ20

  • New shielding and protection will be installed, as well as holes for cabling and equipment.
  • The refurbishment of the painting in the Linac2 was not originally in the planning, so additional budget needs to be allocated.

Reconfiguration of the Beam Emittance Measurement Line (LBE) in the PS SwitchYard

  • The PS switchyard to PSB shielding has been reduced by 2 meters and is awaiting an official assessment from HSE-RP.
  • The beam stoppers will be removed in the LBI line, so a temporary configuration will be used for the beam tests in 2019.

Linac4 beam tests

  • For the individual tests of equipment, a network separation will be needed to allow to power only the transfer line and not the rest of the switchyard.

J. Coupard commented that the schedule is continuous, with little margin. The 3D integration is complete and all documentation and differential drawings are being finalized.

She thanked A. Berjillos for his work as a fellow preparing the Linac4 connection activities.

Comments

  • G. P. Di Giovanni asked about the recent request of POPs commissioning which was mentioned at LS2C to stop the LBE run in 2020 for 3 weeks. J. Coupard answered that the magnet tests need to be done before the closing of the switchyard and after the POPs commissioning in September, which will last two weeks. B. Mikulec asked if the commissioning could be moved forward. J. Coupard answered that it depends on TE-EPC. B. Mikulec suggested keeping one week reserved.
  • G. P. Di Giovanni asked about the origin of the integration drawing budget as there is no explicit budget in the current baseline. J. Coupard suggested asking Y. Muttoni.

AOB

Update on the ECRs (T. Birtwistle)

Multiple groups

  • Magnet modifications in BT, BTM and BTP (TE-MSC, A. Newborough): the integration issues were resolved, waiting for approval.

TE-ABT

  • Replace kicker BT.KFA10 (L. Sermeus, W. Weterings): approved, but the integration model needs to be updated.

TE-MSC

  • Upgrade of the BHZ Main unit bending magnets (A. Newborough): under approval. The cabling issues are now resolved. T. Birtwistle will present the integration drawings next Friday at the IEFC for approval

TE-EPC

  • Upgrade of PS Booster Corrector and Quadrupole Power Supplies (L. D.  Mallac): under approval (integration needed).
  • Booster Power House (BHP) Converters Upgrade (S. Pittet): under approval (integration needed).
  • Installation of the power converters for BI.SMV10 magnets for LIU PSB injection (J-M. Cravero): under approval (integration needed).
  • Installation of new Power Converters for PSB injection Qstrip, Correctors, Quad. (S. Pittet): under approval (integration needed).
  • An estimated planning has been drafted.

BE-BI

  • New BLMs Installation in the LT, LTB and BI Lines for H- exchange monitoring (C. Zamantzas): needed.
  • Upgrade and consolidation of BT.BTV20 (S. Burger): under approval. There are a few points to resolve. M. Hourican identified that the BTV would need to be removed in order to remove the extraction septum – the supports of the BTV will be modified to allow the septa exchange.
  • B. Mikulec mentioned that the ECRs about the renovation of the screens has been rejected by M. Hourican, so this ECR should be closely followed up.

EN-STI

  • New PS Booster Absorber in Period 8 (I. Lamas Garcia): new version sent for approval.

EN-CV

  • Consolidation of the Booster chilled water plant (G. Petrika): under preparation. Integration being finalised.
  • Consolidation of the Booster cooling plant (G. Petrika): under preparation. Integration being finalised.

EN-ACE

  • New LBE line plus infrastructure (J. Coupard): under approval (integration). An updated version will be provided this year.

 

 

There are minutes attached to this event. Show them.